OpenStudy (anonymous):

Can someone check my answer for this? What is the side length b in the triangle below?

OpenStudy (anonymous):

12^2+b^2=20^2 b=16

OpenStudy (vishweshshrimali5):

yes u r correct

OpenStudy (vishweshshrimali5):

gud work

OpenStudy (anonymous):

thank you!! :D

OpenStudy (vishweshshrimali5):

:D